OpenStudy (anonymous):

find the derivative of the function f(x)=(3x-2)^{-3/2} + (x ^{2} +1)^{3/2}

OpenStudy (zehanz):

So\[f(x)=(3x-2)^{-\frac{ 3 }{ 2 }} + (x ^{2} +1)^{\frac{3}{2}}\]Differentiate this function term by term (act as if it were the sum of two separate functions). You can use the Power Rule:\[(x^n)' =nx^{n-1}\]Because there are functions between the brackets, you need to appy the Chain Rule as well. Here, this means multiply the first part with 3 (derivative of 3x-2) and the secont part with 2x (derivative of x²+1).

OpenStudy (zehanz):

If you don't succeed, I can help you out!

OpenStudy (anonymous):

im lost

OpenStudy (zehanz):

OK, first step:\[\left( (3x-2)^{-\frac{3}{2}}\right)'=-\frac{ 3 }{ 2 }(3x-2)^{-\frac{5}{2}} \cdot 3\]See? I used the Power Rule. The -3/2 is now in front of (3x-2) and the exponent is 1 less, -5/2. Also I multiplied with 3, because that is the derivative of 3x-2 (Chain Rule). Now you do the secont one!